Can you please show me how to write an hyperbola equation with foci (o,-2) and (0,2) and vertices of (0,-1) and (0,1).
Answer by lwsshak3(11628)

Can you please show me how to write an hyperbola equation with foci (o,-2) and (0,2) and vertices of (0,-1) and (0,1)
**
Given hyperbola has a vertical transverse axis.
Its standard form of equation: (y-k)^2/a^2-(x-h)^2/b^2=1, (h,k)=(x,y) coordinates of center.
For given hyperbola:
center: (0,0)
length of vertical transverse axis=2 (-1 to 1)=2a
a=1
a^2=1
..
From foci:
2c=4 (-2 to 2)
c=2
c^2=4
..
c^2=a2+b^2
b^2=c^2-a^2=4-1=3
..
Equation of given hyperbola:
y^2/1-x^2/3=1